# The norms for symmetric and antisymmetric tensor products of the weighted shift operators

### v1

@@ Abstract
if and only if the weight satisfies the regularity condition, we partially solve [6, Problem 6 and Problem 7]. It will be seen that most weighted shift operators on function spaces, including weighted Bergman shift, Hardy shift, Dirichlet shift, etc, satisfy the regularity condition. Moreover, at the end of the paper, we solve [6, Problem 1 and Problem 2].
00footnotetext: 2020 AMS Subject Classification: primary 47A13; secondary 46H25.
@@ 1 Introduction
Symmetric tensor products and antisymmetric tensor products are essential in various fields. For instance, symmetric tensors are foundational to general relativity [2]. Additionally, symmetric tensor products play significant roles in multilinear algebra [7], representation theory [5] and statistics [13]. Moreover, decomposing a symmetric tensor into a minimal linear combination of tensor powers of the same vector arises in mobile communications, machine learning, factor analysis of $k$-way arrays, biomedical engineering, psychometrics, and chemometrics [3, 4, 11, 16, 17]. Furthermore, the symmetric part of a quantum geometric tensor has been utilized as a tool to detect quantum phase transitions in PT-symmetric quantum mechanics [18]. In particular, symmetric tensor products of operators are crucial in the study of the many-body system [10] and are actively explored within the physics community [8, 9, 12].
In the present paper, we will study the norm of symmetric and antisymmetric tensor product of operators. Let $\mathcal{H}$ be a Hilbert space and $\mathcal{B}(\mathcal{H})$ be the space of bounded linear operators on $\mathcal{H}$. Let $\mathcal{H}^{\otimes n}$ be the full tensor space of $\mathcal{H}$, $\mathcal{H}^{\odot n}$ be the symmetric tensor space of $\mathcal{H}$ and $\mathcal{H}^{\wedge n}$ be the antisymmetric tensor space of $\mathcal{H}$, as introduced in [14, Page 106]. For $\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{n}\in\mathcal{H}$, let $\mathbf{v}_{1}\odot\mathbf{v}_{2}\odot\cdots\odot\mathbf{v}_{n}$ be the symmetric tensor and $\mathbf{v}_{1}\wedge\mathbf{v}_{2}\wedge\cdots\wedge\mathbf{v}_{n}$ be the antisymmetric tensor of $\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{n}$. For $\mathbb{A}_{1},\mathbb{A}_{2},\cdots,\mathbb{A}_{n}\in\mathcal{B}(\mathcal{H})$, we define $\mathbb{A}_{1}\otimes\mathbb{A}_{2}\otimes\cdots\otimes\mathbb{A}_{n}$ on $\mathcal{H}^{\otimes n}$ as
@@
It is trivial to see that
@@
However the norms of symmetric products and antisymmetric products of bounded operators are much more complicated. For example, in [6, Proposition 7.2], Garcia, O’Loughlin and Yu estimated the norms of symmetric product of diagonal operators $\mathbb{A}_{1}$ and $\mathbb{A}_{2}$, and the sharpness of the following estimates was obtained:
@@
moreover, they raised a series of problems aimed at exploring the norms and spectrums of symmetric and antisymmetric product of operators. To continue, let $\{e_{i}\}_{i=0}^{\infty}$ be an orthonormal basis(ONB for short) of a complex Hilbert space $\cal H$. In what follows, we always assume that the weight $\alpha=\{\alpha_{i}\}_{i=0}^{\infty}$ is convergent. Let $S_{\alpha}$ be a weighted shift operator on $\cal H$ with the weight $\alpha$, defined by
$S_{\alpha}e_{i}=\alpha_{i}e_{i+1}\ \mathrm{for}\ i\geq 0,$
which was deeply studied in [15]. In particular, $S$ is the forward shift operator with the constant weight $\alpha_{i}=1$, for $i\in\mathbb{N}$. The following problems were raised in [6].
Problem 1.1.
([6, Problem 6]) Identify the norm and spectrum of arbitrary symmetric or antisymmetric tensor products of $S$ and $S^{*}$(for example, consider $S^{2}\odot S\odot S^{*3}$ and $S^{2}\wedge S\wedge S^{*3}$).
@@
([6, Problem 7]) Describe the norm and spectrum of $S_{\alpha}\odot S_{\alpha}^{*}$ and $S_{\alpha}\wedge S_{\alpha}^{*}$, in which $S_{\alpha}$ is a weighted shift operator. What can be said if more factors are included?
In this paper, we partially solve these problems. To clarify our main results, we need the assumption of the so-called regularity on the weight.
Definition 1.1.
@@
The following theorem is the main result in the present paper, which partially gives an affirmative answer to the [6, Problem 6 and Problem 7].
@@
    As a consequence of the above theorem, for $(l_{1},\cdots,l_{n})\in\mathbb{N}^{n}$
@@
    which partially solves Problem 1.1.
@@
    It is evident that the shift operators on nearly every classical function space, including the Hardy shift, weighted Bergman shift, and Dirichlet shift, are regular.
This paper is organized as follows. In Section 2, we provide some preparations to facilitate the proof of Theorem 1.3. Section 3 is devoted to the proof of our main result, Theorem 1.3. Finally, in Section 4, we solve [6, Problem 1 and Problem 2].
@@ 2 Preliminaries
Denote $\mathbb{N}^{n}/\Sigma_{n}$ and $\mathbb{Z}^{n}/\Sigma_{n}$ to be the quotient spaces and $\langle i_{1},\cdots,i_{n}\rangle\in\mathbb{Z}^{n}/\Sigma_{n}$ ($\langle\mathfrak{i}\rangle$ for short) is the coset of $\mathfrak{i}=(i_{1},\cdots,i_{n})$. For $\mathfrak{i}\in\mathbb{Z}^{n}$, $\pi\in\Sigma$, set
@@
and it is easy to see that for any $\pi\in\Sigma$, $e_{\mathfrak{i}}=e_{\mathfrak{i}_{\pi}}$ and write
$e_{\langle\mathfrak{i}\rangle}=e_{\mathfrak{i}}.$ (2.7)
@@
$\left\langle l_{1},l_{2},\cdots,l_{n}\right\rangle=\left\langle\underbrace{{\tilde{l}_{1}},\cdots,{\tilde{l}_{1}}}_{n_{1}},\cdots,\underbrace{{\tilde{l}_{k}},\cdots,{\tilde{l}_{k}}}_{n_{k}}\right\rangle,\quad\mathrm{where}\ {\tilde{l}_{i}}\neq{\tilde{l}_{j}}\text{ for }i\not=j,$ (2.8)
@@
$R_{\mathfrak{i}}=\left\{\langle\mathfrak{j}\rangle\in\mathbb{N}^{n}/\Sigma_{n}:\exists\,\pi\in\Sigma_{n}~{}s.t.~{}\langle\mathfrak{j}+\mathfrak{l}_{\pi}\rangle=\langle\mathfrak{i}\rangle\right\},$ (2.10)
@@
#$^{\#}R_{\mathfrak{i}}\leq\ ^{\#}\left\{\mathfrak{l}_{\pi}:\pi\in\Sigma_{n}\ \right\}=\mathscr{M}.$ (2.11)
@@
$A_{k}=\begin{cases}\left\{\langle\mathfrak{i}\rangle\in\mathbb{N}^{n}/\Sigma_{n}:|\mathfrak{i}|=k\ and\ ^{\#}R_{\mathfrak{i}}<\mathscr{M}\right\},&k\geq 0,\\ \emptyset,&k<0.\end{cases}$ (2.12)
@@
Lemma 2.2.
For $k>0,$ if $l_{1},l_{2},\cdots,l_{n}$ are not all equal, then there exists a constant $M>0$, such that ${}^{\#}A_{k}\leq Mk^{n-2}$.
@@
[…]
@@
[…]
@@
${}^{\#}C_{k,j}\leq^{\#}\bigcup\limits_{i_{j}=0}^{h}\left\{\mathfrak{i}\in\mathbb{N}^{n}:|\mathfrak{i}|=k\ \right\}\leq hk^{n-2}.$
Notice that $B_{k}=\bigcup\limits_{j=1}^{n}C_{k,j},$ hence ${}^{\#}B_{k}\leq nhk^{n-2}$. Next we will prove (2.16). For $\mathfrak{i}\in D_{k}$, by the definitions of $D_{k}$ and $A_{k}^{\prime}$, we have
@@
#{$^{\#}\left\{\left\langle\mathfrak{i}-\mathfrak{l}_{\sigma}\right\rangle:\sigma\in\Sigma_{n}\right\}=^{\#}R_{\mathfrak{i}}<\mathscr{M}.$ (2.19)
@@
$\left\langle\mathfrak{i}-\mathfrak{l}_{\tau_{0}}\right\rangle=\left\langle\mathfrak{i}-\mathfrak{l}_{\sigma_{0}}\right\rangle.$ (2.20)
Otherwise, if for all $\sigma,\tau\in\Sigma_{n}$ such that $\mathfrak{l}_{\sigma_{0}}\not=\mathfrak{l}_{\tau_{0}}$, we have
$\left\langle\mathfrak{i}-\mathfrak{l}_{\tau_{0}}\right\rangle\not=\left\langle\mathfrak{i}-\mathfrak{l}_{\sigma_{0}}\right\rangle.$
@@
{${}^{\#}\left\{\left\langle\mathfrak{i}-\mathfrak{l}_{\sigma}\right\rangle:\sigma\in\Sigma_{n}\right\}=\mathscr{M},$
@@
$D_{k,\tau,\sigma}$ $=$ $\left\{\mathfrak{i}\in D_{k}:\left\langle\mathfrak{i}-\mathfrak{l}_{\tau}\right\rangle=\left\langle\mathfrak{i}-\mathfrak{l}_{\sigma}\right\rangle\right\}$
then by (2.20)
@@
{${}^{\#}\left\{\mathfrak{i}\in\mathbb{N}^{n}:|\mathfrak{i}|=k,\ i_{a}-l_{\tau(a)}=i_{b}-l_{\sigma(b)}\ \right\}\leq k^{n-2}.$
@@
$E_{r,m}=\begin{cases}\left\{\left\langle\mathfrak{i}\right\rangle\in\mathbb{N}^{n}/\Sigma_{n}:\left|\mathfrak{i}\right|=r\ and\ ^{\#}R_{\mathfrak{i}}=m\right\},&r\geq 0,\\ \emptyset,&r<0.\end{cases}$ (2.22)
@@
$N_{\mathfrak{j},\mathfrak{i}}=\left\{\pi\in\Sigma_{n}:\left\langle\mathfrak{i}+\mathfrak{l}_{\pi}\right\rangle=\left\langle\mathfrak{j}\right\rangle\right\}.$ (2.23)
@@
#{$^{\#}\left\{\left\langle\mathfrak{j}-\mathfrak{l}_{\pi}\right\rangle:\pi\in\Sigma_{n}\right\}\leq\ ^{\#}\left\{\mathfrak{l}_{\pi}:\pi\in\Sigma_{n}\right\}-1.$ (2.24)
@@
Define $T\colon\,\left\{\mathfrak{l}_{\pi}:\pi\in\Sigma_{n}\right\}\to\left\{\left\langle\mathfrak{j}-\mathfrak{l}_{\pi}\right\rangle:\pi\in\Sigma_{n}\right\}$ by
$T:\mathfrak{l}_{\pi}\mapsto\left\langle\mathfrak{j}-\mathfrak{l}_{\pi}\right\rangle.$
@@
$T\left(\mathfrak{l}_{\pi_{1}}\right)=\left\langle\mathfrak{j}-\mathfrak{l}_{\pi_{1}}\right\rangle=\left\langle\mathfrak{j}-\mathfrak{l}_{\pi_{2}}\right\rangle=T\left(\mathfrak{l}_{\pi_{2}}\right),$
@@
$T\left(\mathfrak{l}_{\pi_{1}}\right)=\left\langle\mathfrak{j}-\mathfrak{l}_{\pi_{1}}\right\rangle=\left\langle\mathfrak{j}-\mathfrak{l}_{\pi_{2}}\right\rangle=T\left(\mathfrak{l}_{\pi_{2}}\right),$
which implies that $T$ is not injective.
@@
Suppose that $l_{1},l_{2},\cdots,l_{n}$ are not all equal. Then for $k>0$ large enough, and every $\mathfrak{i},\mathfrak{j}\in\mathbb{N}^{n},$ which satisfies that $\langle\mathfrak{j}\rangle\in E_{k+S_{\mathfrak{l}},\mathscr{M}}$ and $\left\langle\mathfrak{i}\right\rangle\in R_{\mathfrak{j}},$ we have
$\left\|e_{\langle\mathfrak{i}\rangle}\right\|\leq\frac{{}^{\#}N_{\mathfrak{j},\mathfrak{i}}\mathscr{M}}{n!\sqrt{n!}}$
@@
Since $\langle\mathfrak{j}\rangle\in E_{k+S_{\mathfrak{l}},\mathscr{M}}$, by the definition of $E_{k+S_{\mathfrak{l}},\mathscr{M}}$ in (2.22),
#$^{\#}R_{\mathfrak{j}}=\mathscr{M}.$ (2.25)
@@
$R_{\mathfrak{j}}=\left\{\left\langle\mathfrak{j}-\mathfrak{l}_{\pi}\right\rangle:\pi\in\Sigma_{n},j_{i}-l_{\pi(i)}\geq 0\right\}.$
@@
${}^{\#}R_{\mathfrak{j}}\leq\ ^{\#}\left\{\left\langle\mathfrak{j}-\mathfrak{l}_{\pi}\right\rangle:\pi\in\Sigma_{n}\right\}\leq\ ^{\#}\left\{\mathfrak{l}_{\pi}:\pi\in\Sigma_{n}\right\}-1=\ \mathscr{M}-1,$
@@
$\left\langle i_{1},\cdots,i_{n}\right\rangle=\left\langle\underbrace{\widetilde{i_{1}},\cdots,\widetilde{i_{1}}}_{t_{1}},\cdots,\underbrace{\widetilde{i_{s}},\cdots,\widetilde{i_{s}}}_{t_{s}}\right\rangle,\mathrm{where}\ \widetilde{i_{1}}\neq\cdots\neq\widetilde{i_{s}}.$
Since $\left\langle\mathfrak{i}\right\rangle\in R_{\mathfrak{j}},$ we have $N_{\mathfrak{j},\mathfrak{i}}\neq\emptyset.$ Therefore by (2.23) and (2.26),
${}^{\#}N_{\mathfrak{j},\mathfrak{i}}$ =#$=\ ^{\#}\left\{\pi\in\Sigma_{n}:\left\langle\mathfrak{i}+\mathfrak{l}_{\pi}\right\rangle=\left\langle\mathfrak{j}\right\rangle\right\}\geq\ t_{1}!t_{2}!\cdots t_{s}!n_{1}!\cdots n_{k}!.$
@@
$\left\|e_{\langle\mathfrak{i}\rangle}\right\|$ $=$ $\sqrt{\frac{t_{1}!t_{2}!\cdots t_{s}!}{n!}}\leq\frac{t_{1}!t_{2}!\cdots t_{s}!}{\sqrt{n!}}\leq\frac{{}^{\#}N_{\mathfrak{j},\mathfrak{i}}}{n_{1}!n_{2}!\cdots n_{k}!\sqrt{n!}}=\frac{{}^{\#}N_{\mathfrak{j},\mathfrak{i}}\mathscr{M}}{n!\sqrt{n!}}.$
@@
$P(k,n)=\ ^{\#}\left\{\left\langle k_{1},\cdots,k_{n}\right\rangle:\sum\limits_{i=1}^{n}k_{i}=k,k_{i}\geq 0\right\}.$ (2.28)
Moreover, $Q(k,n)$ is defined by the number of partitions of $k$ into at most $n$ ordered parts, i.e.
$Q(k,n)=\ ^{\#}\left\{\left(k_{1},\cdots,k_{n}\right):\sum\limits_{i=1}^{n}k_{i}=k,k_{i}\geq 0\right\}.$
@@
$\sum\limits_{k=0}^{\infty}P(k,m)q^{n}=\frac{1}{\left(1-q\right)\left(1-q^{2}\right)\cdots\left(1-q^{m}\right)}.$
@@
$\sum\limits_{k=0}^{\infty}P(k,m+1)q^{n}$ $=\left(\sum\limits_{k=0}^{\infty}P(k,m)q^{n}\right)\frac{1}{1-q^{m+1}}$
$=\left(\sum\limits_{k=0}^{\infty}P(k,m)q^{n}\right)\sum\limits_{k=0}^{\infty}\left(q^{m+1}\right)^{k}$
@@
$=\sum\limits_{i=0}^{s}P(\left(m+1\right)i+t+1,m).$
@@
$\lim_{s\to\infty}\frac{P\left(\left(m+1\right)s+t,m+1\right)}{P\left(\left(m+1\right)s+t+1,m+1\right)}$ $=\lim_{s\to\infty}\frac{\sum\limits_{i=0}^{s}P(\left(m+1\right)i+t,m)}{\sum\limits_{i=0}^{s}P(\left(m+1\right)i+t+1,m)}$
@@
$=1.$
For $01>\varepsilon>$ 0, let
$\tilde{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}=\left\{\left\langle\mathfrak{j}\right\rangle\in E_{k+S_{\mathfrak{l}},\mathscr{M}}:\Gamma_{j_{m},-l_{\pi(m)}}>1-\varepsilon,\forall 1\leq m\leq n,\forall\,\pi\in\Sigma_{n}\right\}$ (2.33)

### v2

@@ Abstract
if and only if the weight satisfies the regularity condition, we partially solve [7, Problem 6 and Problem 7]. It will be seen that many weighted shift operators on function spaces, including weighted Bergman shift, Hardy shift, etc, satisfy the regularity condition. Moreover, at the end of the paper, we solve [7, Problem 1 and Problem 2].
00footnotetext: 2020 AMS Subject Classification: primary 47A30; secondary 47A80.
@@ 1 Introduction
Symmetric tensor products and antisymmetric tensor products are essential in various fields. For instance, symmetric tensors are foundational to general relativity [3]. Additionally, symmetric tensor products play significant roles in multilinear algebra [8], representation theory [6] and statistics [15]. Moreover, decomposing a symmetric tensor into a minimal linear combination of tensor powers of the same vector arises in mobile communications, machine learning, factor analysis of $k$-way arrays, biomedical engineering, psychometrics, and chemometrics [4, 5, 13, 18, 19]. Furthermore, the symmetric part of a quantum geometric tensor has been utilized as a tool to detect quantum phase transitions in PT-symmetric quantum mechanics [21]. In particular, symmetric tensor products of operators are crucial in the study of non-interacting systems [12] and are actively explored within the physics community [9, 10, 14].
In the present paper, we will study the norm of symmetric and antisymmetric tensor product of operators. Let $\mathcal{H}$ be a Hilbert space and $\mathcal{B}(\mathcal{H})$ be the space of bounded linear operators on $\mathcal{H}$. Let $\mathcal{H}^{\otimes n}$ be the full tensor space of $\mathcal{H}$, $\mathcal{H}^{\odot n}$ be the symmetric tensor space of $\mathcal{H}$ and $\mathcal{H}^{\wedge n}$ be the antisymmetric tensor space of $\mathcal{H}$, as introduced in [16, Page 106]. For $\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{n}\in\mathcal{H}$, let $\mathbf{v}_{1}\odot\mathbf{v}_{2}\odot\cdots\odot\mathbf{v}_{n}$ be the symmetric tensor and $\mathbf{v}_{1}\wedge\mathbf{v}_{2}\wedge\cdots\wedge\mathbf{v}_{n}$ be the antisymmetric tensor of $\mathbf{v}_{1},\mathbf{v}_{2},\cdots,\mathbf{v}_{n}$. For $\mathbb{A}_{1},\mathbb{A}_{2},\cdots,\mathbb{A}_{n}\in\mathcal{B}(\mathcal{H})$, we define $\mathbb{A}_{1}\otimes\mathbb{A}_{2}\otimes\cdots\otimes\mathbb{A}_{n}$ on $\mathcal{H}^{\otimes n}$ as
@@
By [2, Page 162] and [11, (3.8.17)],
$\sigma(\mathbb{A}_{1}\otimes\mathbb{A}_{2}\otimes\cdots\otimes\mathbb{A}_{n})=\sigma(\mathbb{A}_{1})\cdots\sigma(\mathbb{A}_{n})$
and
@@
However the spectra and norms of symmetric tensor products and antisymmetric tensor products of bounded operators are much more complicated. For example, in [7, Proposition 7.2], Garcia, O’Loughlin and Yu estimated the norms of symmetric tensor product of diagonal operators $\mathbb{A}_{1}$ and $\mathbb{A}_{2}$, and the sharpness of the following estimates was obtained:
@@
[…]
$S_{\alpha}e_{i}=\alpha_{i}e_{i+1}\quad\text{for }i\geq 0,$
which has been extensively studied in [17]. Let $S_{\alpha}^{*}$ denote its adjoint operator. In particular, $S$ is the forward shift operator with the constant weight $\alpha_{i}=1$ for $i\in\mathbb{N}$.
The following problems were raised in [7].
Problem $1.$
([7, Problem 6]) Identify the norm and spectrum of arbitrary symmetric or antisymmetric tensor products of $S$ and $S^{*}$(for example, consider $S^{2}\odot S\odot S^{*3}$ and $S^{2}\wedge S\wedge S^{*3}$).
@@
([7, Problem 7]) Describe the norm and spectrum of $S_{\alpha}\odot S_{\alpha}^{*}$ and $S_{\alpha}\wedge S_{\alpha}^{*}$, in which $S_{\alpha}$ is a weighted shift operator. What can be said if more factors are included?
We note that Yang and Zhang [20] addressed spectral-related aspects of Problem 1.2 very recently. In this paper, we partially solve these problems. To clarify our main results, we need the assumption of the so-called regularity on the weight.
Definition $1.$
@@
The following theorem is the main result in the present paper, which partially gives an affirmative answer to Problem 1.1 and Problem 1.2.
@@
    As a consequence of the above theorem, for $(l_{1},\cdots,l_{n})\in\mathbb{N}^{n},$
@@
    which partially solves Problem $1.$
@@
    It is evident that the shift operators on nearly every classical function space, including the Hardy shift and weighted Bergman shift, are regular.
This paper is organized as follows. In Section 2, we provide some preparations to facilitate the proof of Theorem 1.3. Section 3 is devoted to the proof of our main result, Theorem 1.3. Finally, in Section 4, we solve [7, Problem 1 and Problem 2].
@@ 2 Preliminaries
Denote $\mathbb{N}^{n}/\Sigma_{n}$ and $\mathbb{Z}^{n}/\Sigma_{n}$ to be the quotient spaces and $i_{1},\cdots,i_{n}]\in\mathbb{Z}^{n}/\Sigma_{n}$ ($\mathfrak{i}]$ for short) is the coset of $\mathfrak{i}=(i_{1},\cdots,i_{n})$. For $\mathfrak{i}\in\mathbb{Z}^{n}$, $\pi\in\Sigma_{n}$, set
@@
and it is easy to see that for any $\pi\in\Sigma_{n}$, $e_{\mathfrak{i}}=e_{\mathfrak{i}_{\pi}}$ and write
$e_{[\mathfrak{i}]}=e_{\mathfrak{i}}.$ (2.7)
@@
$[l_{1},l_{2},\cdots,l_{n}]=[\underbrace{{\tilde{l}_{1}},\cdots,{\tilde{l}_{1}}}_{n_{1}},\cdots,\underbrace{{\tilde{l}_{k}},\cdots,{\tilde{l}_{k}}}_{n_{k}}],\quad\mathrm{where}\ {\tilde{l}_{i}}\neq{\tilde{l}_{j}}\text{ for }i\not=j,$ (2.8)
@@
$R_{\mathfrak{i}}=\left\{[\mathfrak{j}]\in\mathbb{N}^{n}/\Sigma_{n}:\exists\,\pi\in\Sigma_{n}~s.t.~[\mathfrak{j}+\mathfrak{l}_{\pi}]=[\mathfrak{i}]\right\},$ (2.10)
@@
$^{\#}R_{\mathfrak{i}}\leq\ ^{\#}\left\{\mathfrak{l}_{\pi}:\pi\in\Sigma_{n}\ \right\}=\mathscr{M}.$ (2.11)
@@
$A_{k}=\begin{cases}\left\{[\mathfrak{i}]\in\mathbb{N}^{n}/\Sigma_{n}:|\mathfrak{i}|=k\ and\ ^{\#}R_{\mathfrak{i}}<\mathscr{M}\right\},&k\geq 0,\\ \emptyset,&k<0.\end{cases}$ (2.12)
@@
Lemma $2.$
For $k>0$ sufficiently large, if $l_{1},l_{2},\cdots,l_{n}$ are not all equal, then there exists a constant $M>0$, such that $^{\#}A_{k}\leq Mk^{n-2}$.
@@
$B_{k}=\left\{\mathfrak{i}\in A_{k}^{\prime}:\exists\,1\leq j\leq n\ and\ \exists\,\pi\in\Sigma_{n}\ such\ that\ i_{j}-l_{\pi(j)}<0\ \right\},$
@@
$D_{k}=\left\{\mathfrak{i}\in A_{k}^{\prime}:\forall\,1\leq j\leq n\ and\ \forall\,\pi\in\Sigma_{n}\ such\ that\ i_{j}-l_{\pi(j)}\geq 0\ \right\}.$
@@
${}^{\#}C_{k,j}\leq^{\#}\bigcup\limits_{i_{j}=0}^{h}\left\{\mathfrak{i}\in\mathbb{N}^{n}:|\mathfrak{i}|=k\ \right\}\leq\left(h+1\right)\left(k+1\right)^{n-2}.$
Notice that $B_{k}=\bigcup\limits_{j=1}^{n}C_{k,j},$ hence $^{\#}B_{k}\leq n\left(h+1\right)\left(k+1\right)^{n-2}$. Next we will prove (2.16). For $\mathfrak{i}\in D_{k}$, by the definitions of $D_{k}$ and $A_{k}^{\prime}$, we have
@@
$^{\#}\left\{[\mathfrak{i}-\mathfrak{l}_{\sigma}]:\sigma\in\Sigma_{n}\right\}=^{\#}R_{\mathfrak{i}}<\mathscr{M}.$ (2.19)
@@
$\mathfrak{i}-\mathfrak{l}_{\tau_{0}}]=[\mathfrak{i}-\mathfrak{l}_{\sigma_{0}}].$ (2.20)
Otherwise, if for all $\sigma,\tau\in\Sigma_{n}$ such that $\mathfrak{l}_{\sigma}\not=\mathfrak{l}_{\tau}$, we have
$[\mathfrak{i}-\mathfrak{l}_{\tau}]\not=[\mathfrak{i}-\mathfrak{l}_{\sigma}].$
@@
$^{\#}\left\{[\mathfrak{i}-\mathfrak{l}_{\sigma}]:\sigma\in\Sigma_{n}\right\}$$\mathscr{M}$,
@@
$D_{k,\tau,\sigma}$ $=$ $\left\{\mathfrak{i}\in D_{k}:\exists 1\leq a\neq b\leq n~s.t.~i_{a}-l_{\tau(a)}=i_{b}-l_{\sigma(b)}\right\}.$
Obviously,
@@
$^{\#}\left\{\mathfrak{i}\in\mathbb{N}^{n}:|\mathfrak{i}|=k,\ i_{a}-l_{\tau(a)}=i_{b}-l_{\sigma(b)}\ \right\}\leq\left(k+1\right)^{n-2}.$
@@
$E_{r,m}=\begin{cases}\left\{[\mathfrak{i}]\in\mathbb{N}^{n}/\Sigma_{n}:\left|\mathfrak{i}\right|=r\ and\ ^{\#}R_{\mathfrak{i}}=m\right\},&r\geq 0,\\ \emptyset,&r<0.\end{cases}$ (2.22)
@@
$N_{\mathfrak{j},\mathfrak{i}}=\left\{\pi\in\Sigma_{n}:[\mathfrak{i}+\mathfrak{l}_{\pi}]=[\mathfrak{j}]\right\}.$ (2.23)
@@
$^{\#}\left\{[\mathfrak{j}-\mathfrak{l}_{\pi}]:\pi\in\Sigma_{n}\right\}\leq\ ^{\#}\left\{\mathfrak{l}_{\pi}:\pi\in\Sigma_{n}\right\}-1.$ (2.24)
@@
Define $T\colon\,\left\{\mathfrak{l}_{\pi}:\pi\in\Sigma_{n}\right\}\to\left\{[\mathfrak{j}-\mathfrak{l}_{\pi}]:\pi\in\Sigma_{n}\right\}$ by
$T:\mathfrak{l}_{\pi}\mapsto[\mathfrak{j}-\mathfrak{l}_{\pi}].$
@@
$T\left(\mathfrak{l}_{\pi_{1}}\right)=[\mathfrak{j}-\mathfrak{l}_{\pi_{1}}]=[\mathfrak{j}-\mathfrak{l}_{\pi_{2}}]=T\left(\mathfrak{l}_{\pi_{2}}\right),$
@@
$T\left(\mathfrak{l}_{\pi_{1}}\right)=[\mathfrak{j}-\mathfrak{l}_{\pi_{1}}]=[\mathfrak{j}-\mathfrak{l}_{\pi_{2}}]=T\left(\mathfrak{l}_{\pi_{2}}\right),$
which implies that $T$ is not injective. ∎
@@
Suppose that $l_{1},l_{2},\cdots,l_{n}$ are not all equal. Then for $k>0$ large enough, and every $\mathfrak{i},\mathfrak{j}\in\mathbb{N}^{n},$ which satisfies that $\mathfrak{j}]\in E_{k+S_{\mathfrak{l}},\mathscr{M}}$ and $\mathfrak{i}]\in R_{\mathfrak{j}},$ we have
$\left\|e_{[\mathfrak{i}]}\right\|\leq\frac{{}^{\#}N_{\mathfrak{j},\mathfrak{i}}\mathscr{M}}{n!\sqrt{n!}}.$
@@
Since $\mathfrak{j}]\in E_{k+S_{\mathfrak{l}},\mathscr{M}}$, by the definition of $E_{k+S_{\mathfrak{l}},\mathscr{M}}$ in (2.22),
$^{\#}R_{\mathfrak{j}}=\mathscr{M}.$ (2.25)
@@
$R_{\mathfrak{j}}=\left\{[\mathfrak{j}-\mathfrak{l}_{\pi}]:\pi\in\Sigma_{n},j_{i}-l_{\pi(i)}\geq 0,\forall\,1\leq i\leq n\right\}.$
@@
${}^{\#}R_{\mathfrak{j}}\leq\ ^{\#}\left\{[\mathfrak{j}-\mathfrak{l}_{\pi}]:\pi\in\Sigma_{n}\right\}\leq\ ^{\#}\left\{\mathfrak{l}_{\pi}:\pi\in\Sigma_{n}\right\}-1=\ \mathscr{M}-1,$
@@
$i_{1},\cdots,i_{n}]=[\underbrace{\widetilde{i_{1}},\cdots,\widetilde{i_{1}}}_{t_{1}},\cdots,\underbrace{\widetilde{i_{s}},\cdots,\widetilde{i_{s}}}_{t_{s}}],\mathrm{where}\ \widetilde{i_{1}}\neq\cdots\neq\widetilde{i_{s}}$.
Since $\mathfrak{i}]\in R_{\mathfrak{j}},$ we have $N_{\mathfrak{j},\mathfrak{i}}\neq\emptyset.$ Therefore by (2.23) and (2.26),
${}^{\#}N_{\mathfrak{j},\mathfrak{i}}$ $=\ ^{\#}\left\{\pi\in\Sigma_{n}:[\mathfrak{i}+\mathfrak{l}_{\pi}]=[\mathfrak{j}]\right\}\geq\ t_{1}!t_{2}!\cdots t_{s}!n_{1}!\cdots n_{k}!.$
@@
$\left\|e_{[\mathfrak{i}]}\right\|=\sqrt{\frac{t_{1}!t_{2}!\cdots t_{s}!}{n!}}\leq\frac{t_{1}!t_{2}!\cdots t_{s}!}{\sqrt{n!}}\leq\frac{{}^{\#}N_{\mathfrak{j},\mathfrak{i}}}{n_{1}!n_{2}!\cdots n_{k}!\sqrt{n!}}=\frac{{}^{\#}N_{\mathfrak{j},\mathfrak{i}}\mathscr{M}}{n!\sqrt{n!}}.\qed$
@@
[…]
For $k<0$, $n\geq 1$, let $P(k,n)=0$. Moreover, for $k\geq 0$, $Q(k,n)$ is defined by the number of partitions of $k$ into at most $n$ ordered parts, i.e.
[…]
@@
$\sum\limits_{k=0}^{\infty}P(k,m)q^{k}=\frac{1}{\left(1-q\right)\left(1-q^{2}\right)\cdots\left(1-q^{m}\right)}.$
@@
$\sum\limits_{k=0}^{\infty}P(k,m+1)q^{k}$ $=\left(\sum\limits_{k=0}^{\infty}P(k,m)q^{k}\right)\frac{1}{1-q^{m+1}}$
$=\left(\sum\limits_{k=0}^{\infty}P(k,m)q^{k}\right)\sum\limits_{k=0}^{\infty}\left(q^{m+1}\right)^{k}$
@@
$=\sum\limits_{i=-1}^{s}P(\left(m+1\right)i+t+1,m).$
@@
$\lim_{s\to\infty}\frac{P\left(\left(m+1\right)s+t,m+1\right)}{P\left(\left(m+1\right)s+t+1,m+1\right)}$ $=\lim_{s\to\infty}\frac{\sum\limits_{i=0}^{s}P(\left(m+1\right)i+t,m)}{\sum\limits_{i=-1}^{s}P(\left(m+1\right)i+t+1,m)}$
@@
$=1.\qed$
For $1>\varepsilon>0$, let
$\tilde{A}_{k+S_{\mathfrak{l}},\mathscr{M},\varepsilon}=\left\{[\mathfrak{j}]\in E_{k+S_{\mathfrak{l}},\mathscr{M}}:\Gamma_{j_{m},-l_{\pi(m)}}>1-\varepsilon,\forall 1\leq m\leq n,\forall\,\pi\in\Sigma_{n}\right\}$ (2.33)